# Infinite Dimensional Topological-Holomorphic Symmetry in Three-Dimensions

### v2

@@ 7. The Raviolo Vertex Algebra
A central contribution of [22] is the introduction of the raviolo vertex algebra, a three-dimensional analogue of the standard vertex algebra. From the perspective of quantum field theory, it encodes the algebra of local operators in a theory that is partially holomorphic and partially topological, just as ordinary vertex algebras capture the operator algebra of two-dimensional chiral theories.
@@ 7.1. A brief review of raviolo vertex algebras
  • ${\mathcal{V}}=\oplus{\mathcal{V}}^{r}$ is the space of states, which is $\mathbf{Z}$}-graded
@@ 7.2. Raviolo Current Vertex Algebra
In this final section, we construct the raviolo current vertex algebra associated with the anti-chiral sector of the topological-holomorphic theory (4.2.1). Our strategy is to leverage the canonical representation of $\widehat{\mathfrak{G}}_{\mathfrak{w}}$ on its universal envelope, then apply the reconstruction theorem (Prop. 4.0.1 of [22]) introduced above.
Let us begin by constructing the vacuum module $\mathcal{V}$. From a physicist’s perspective, the vacuum module is simply the Fock space — the vector space generated by arbitrary products of creation operators acting on the vacuum. Since our raviolo vertex algebra will be built from the affine graded Lie algebra $\widehat{\mathfrak{G}}_{\mathfrak{w}}$ constructed in §6, we must identify, among its generators, which act as creation and which as annihilation operators.
@@
To formalize this intuitive picture, one can describe the construction of $\mathcal{V}$ in purely algebraic terms. Consider the affine graded Lie algebra $\widehat{\mathfrak{G}}_{\mathfrak{w}}$ generated by the elements $\ell^{a}_{n}=s^{a}\otimes\Omega_{u}^{n}$ in degree 1 and $\eta^{a}_{n}=t^{a}\otimes u^{n}$ in degree 0. Its universal enveloping graded algebra, denoted $U\widehat{\mathfrak{G}}_{\mathfrak{w}}$, is defined as the free graded algebra generated by polynomials in $\ell_{n}^{a}$ and $\eta_{n}^{a}$, together with a central element $\mathfrak{w}$ of degree 1.
@@
the mode generators ${\tilde{L}}^{a}_{n},{\tilde{H}}^{a}_{n}$ act canonically on $U\hat{\mathfrak{G}}_{\mathfrak{w}}$ by multiplication by $\ell^{a}_{n},\eta^{a}_{n}$, respectively. We will use (7.2.3) extensively in the following.
Next, consider the central factor $\mathbb{C}_{w}=\mathbb{C}\oplus\mathbb{C}w[-1]$, where $w$ is an abstract Abelian generator of degree 1, and whose homogeneous degree-0 unit is $1\oplus 0w=1$. Taking the tensor product $U\widehat{\mathfrak{G}}_{\mathfrak{w}}\otimes\mathbb{C}_{w}$, the vacuum module $\mathcal{V}$ is obtained by quotienting out two relations: (i) all of the $\tilde{H}$}-modes act (through $\rho$) trivially on $\mathcal{V}$, and (ii) the generator of the central piece $\mathbb{C}[1]\subset\widehat{\mathfrak{G}}_{\mathfrak{w}}$ acts as multiplication by the constant ww. [^10] Note that both the values of the 2-cocycle $\mathfrak{w}$ and the scalar $w$ have cohomological degree 1. Equivalently,
(7.2.4) $\mathcal{V}=U\widehat{\mathfrak{G}}_{\mathfrak{w}}\otimes_{U\widehat{\mathfrak{G}}_{\geq 0}}\mathbb{C}_{w}\,$,
where $\widehat{\mathfrak{G}}_{\geq 0}=\mathrm{span}\{{\tilde{H}}_{n}^{a}\}\oplus\mathbb{C}w[-1]$ denotes the positive graded subalgebra.
@@
The vacuum module $\mathcal{V}=\operatorname{Ind}_{U\hat{\mathfrak{G}}_{\geq 0}}^{U\hat{\mathfrak{G}}_{\mathfrak{w}}}{\mathbb{C}}_{w}$ is the induced representation of the trivial $\hat{\mathfrak{G}}_{\mathfrak{w}}$-representation on ${\mathbb{C}}_{w}\cong{\mathbb{C}}\oplus{\mathbb{C}}[1]$.
In the following, we shall leverage (7.2.3) in order to distinguish a vector $\ell^{a}_{n},\eta^{a}_{n}\in\widehat{\mathfrak{G}}_{\mathfrak{w}}\subset\mathcal{V}$, and its corresponding operator ${\tilde{L}}^{a}_{n},{\tilde{H}}^{a}_{n}$ on $\mathcal{V}$.
@@
Let us take the vector space $\mathcal{V}$ defined in (7.2.4) and consider, for $1\leq a\leq\operatorname{dim}\mathfrak{n}=\operatorname{dim}V$, the raviolo field operators
(7.2.5) $\mathcal{J}^{a}(z)=\sum_{n\geq 0}z^{n}{\tilde{L}}_{n}^{a}+\Omega_{z}^{n}{\tilde{H}}^{a}_{n}\,$,
which are homogeneous of degree 0. We identify the unit $|0\rangle=1\otimes 1\in\mathcal{V}$ as the vacuum. We will now prove points (1) through (4) of Proposition 7.1.6.
@@
which proves point (1). Next, for the translation operator, we begin by defining an additional generator $D$ on $\widehat{\mathfrak{G}}_{\mathfrak{w}}\cong(\mathfrak{G}^{\vee}\otimes\mathcal{K}^{\bullet}_{u})\oplus{\mathbb{C}}[1]$ from the action of the derivative $\partial_{u}$ on the current modes:
$D,X\otimes k(u)]=-X\otimes\partial_{u}k(u),\qquad[D,\mathfrak{w}]=0\,$,
where $X\in\mathfrak{G}^{\vee},~k(u)\in\mathcal{K}_{u}^{\bullet}$ and $\mathfrak{w}\in{\mathbb{C}}[1]$ is the generator. Under the canonical $\widehat{\mathfrak{G}}_{\mathfrak{w}}$-representation on $\mathcal{V}$, this gives rise to an endomorphism $\partial\in\operatorname{End}\mathcal{V}$ which, by construction, satisfies
@@
for all $x\in\widehat{\mathfrak{G}}_{\mathfrak{w}}$. More explicitly, these read as the following mode relations
@@
By performing a detailed computation that properly incorporates the grading, we verify that the commutator between any two raviolo fields satisfies
(7.2.10) $\mathcal{J}^{a}(z),\mathcal{J}^{b}(w)]=\partial_{w}\Delta(z-w)(\kappa^{-1})^{ab}+\Delta(z-w)\tilde{\textbf{f}}^{ab}_{c}\mathcal{J}^{c}(w)\,$,
for some constants $\tilde{\textbf{f}}^{ab}_{c}$ and thus, by Lemma (7.1.4) we obtain their mutual locality thereby proving point (3).
@@
Finally, point (4) follows inmmediately by construction of $\mathcal{V}$, and thus, by the reconstruction theorem 7.1.6, there exist a unique raviolo vertex algebra structure on $\mathcal{V}$.

### v3

@@ 7. The Raviolo Vertex Algebra
A key ingredient in [22] is the introduction of the raviolo vertex algebra, a three-dimensional analogue of the standard vertex algebra. From the perspective of quantum field theory, it encodes the algebra of local operators in a theory that is partially holomorphic and partially topological, just as ordinary vertex algebras capture the operator algebra of two-dimensional chiral theories.
@@ 7.1. A brief review of raviolo vertex algebras
  • ${\mathcal{V}}=\oplus{\mathcal{V}}^{r}$ is the space of states, which is $\mathbb{Z}$-graded
@@ 7.2. Raviolo Current Vertex Algebra
In this final section, we construct the raviolo current vertex algebra associated with the anti-chiral sector of the topological-holomorphic theory (4.2.1). Our strategy is to leverage the canonical representation of $\widehat{\mathfrak{G}}_{\mathfrak{w}}^{\vee}$ on its universal envelope, then apply the reconstruction theorem (Prop. 4.0.1 of [22]) introduced above.
Let us begin by constructing the vacuum module $\mathcal{V}$. From a physicist’s perspective, the vacuum module is simply the Fock space — the vector space generated by arbitrary products of creation operators acting on the vacuum. Since our raviolo vertex algebra will be built from the affine graded Lie algebra $\widehat{\mathfrak{G}}_{\mathfrak{w}}^{\vee}$ constructed in §6, we must identify, among its generators, which act as creation and which as annihilation operators.
@@
To formalize this intuitive picture, one can describe the construction of $\mathcal{V}$ in purely algebraic terms. Consider the affine graded Lie algebra $\widehat{\mathfrak{G}}^{\vee}_{\mathfrak{w}}$, with the grading conventions (6.1.14), generated by the homogeneous basis elements $\ell^{a}_{n}=s^{a}\otimes\Omega_{u}^{n}$ in degree 1 and $\eta^{a}_{n}=t^{a}\otimes u^{n}$ in degree 0. Its universal enveloping graded algebra, denoted $U\widehat{\mathfrak{G}}_{\mathfrak{w}}^{\vee}$, is defined as the free graded algebra generated by polynomials in $\ell_{n}^{a}$ and $\eta_{n}^{a}$, together with a central element $\mathfrak{w}$ of degree 1.
@@
the mode generators ${\tilde{L}}^{a}_{n},{\tilde{H}}^{a}_{n}$ act canonically on $U\widehat{\mathfrak{G}}_{\mathfrak{w}}^{\vee}$ by multiplication by $\ell^{a}_{n},\eta^{a}_{n}$, respectively. We will use (7.2.3) extensively in the following.
Next, consider the central factor $\mathbb{C}_{w}=\mathbb{C}\oplus\mathbb{C}w[-1]$, where $w$ is an abstract Abelian generator of degree 1, and whose homogeneous degree-0 unit is $1\oplus 0w=1$. Taking the tensor product $U\widehat{\mathfrak{G}}^{\vee}_{\mathfrak{w}}\otimes\mathbb{C}_{w}$, the vacuum module $\mathcal{V}$ is obtained by quotienting out two relations: (i) all of the $\tilde{H}$}-modes act trivially on $\mathcal{V}$, and (ii) the generator of the central piece $\mathbb{C}[-1]\subset\widehat{\mathfrak{G}}^{\vee}_{\mathfrak{w}}$ acts as multiplication by the constant ww. [^11] Note that both the values of the 2-cocycle $\mathfrak{w}$ and the scalar $w$ have cohomological degree 1. Equivalently,
(7.2.4) $\mathcal{V}=U\widehat{\mathfrak{G}}^{\vee}_{\mathfrak{w}}\otimes_{U\widehat{\mathfrak{G}}^{\vee}_{\geq 0}}\mathbb{C}_{w}\,$,
where $\widehat{\mathfrak{G}}^{\vee}_{\geq 0}=\mathrm{span}\{{\tilde{H}}_{n}^{a}\}\oplus\mathbb{C}w[-1]$ denotes the positive graded subalgebra.
@@
The vacuum module $\mathcal{V}=\operatorname{Ind}_{U\widehat{\mathfrak{G}}^{\vee}_{\geq 0}}^{U\widehat{\mathfrak{G}}^{\vee}_{\mathfrak{w}}}{\mathbb{C}}_{w}$ is the induced representation of the trivial $\widehat{\mathfrak{G}}^{\vee}_{\mathfrak{w}}$-representation on ${\mathbb{C}}_{w}\cong{\mathbb{C}}\oplus{\mathbb{C}}[-1]$.
In the following, we shall leverage (7.2.3) in order to distinguish a vector $\ell^{a}_{n},\eta^{a}_{n}\in\widehat{\mathfrak{G}}^{\vee}_{\mathfrak{w}}\subset\mathcal{V}$, and its corresponding operator ${\tilde{L}}^{a}_{n},{\tilde{H}}^{a}_{n}$ on $\mathcal{V}$.
@@
Let us take the vector space $\mathcal{V}$ defined in (7.2.4) and consider, for $1\leq a\leq\operatorname{dim}\mathfrak{n}^{\vee}=\operatorname{dim}V^{\vee}$, the raviolo field operators
(7.2.5) $\mathcal{J}^{a}(z)=\sum_{n\geq 0}z^{n}{\tilde{L}}_{n}^{a}+\Omega_{z}^{n}{\tilde{H}}^{a}_{n}$
which, with the degree conventions in Remark 5, are homogeneous of degree $1$. We identify the unit $|0\rangle=1\otimes 1\in\mathcal{V}$ as the vacuum. We will now prove points (1) through (4) of Proposition 7.1.6.
@@
which proves point (1). Next, for the translation operator, we begin by defining an additional generator $D$ on $\widehat{\mathfrak{G}}^{\vee}_{\mathfrak{w}}\cong(\mathfrak{G}^{\vee}\otimes\mathcal{K}^{\bullet}_{u})\oplus{\mathbb{C}}[-1]$ from the action of the derivative $\partial_{u}$ on the current modes:
$D,X\otimes k(u)]=X\otimes\partial_{u}k(u),\qquad[D,\mathfrak{w}]=0\,$,
where $X\in\mathfrak{G}^{\vee},~k(u)\in\mathcal{K}_{u}^{\bullet}$ and $\mathfrak{w}\in{\mathbb{C}}[-1]$ is the degree 1 generator. Under the canonical $\widehat{\mathfrak{G}}^{\vee}_{\mathfrak{w}}$-representation on $\mathcal{V}$, this gives rise to an endomorphism $\partial\in\operatorname{End}\mathcal{V}$ which, by construction, satisfies
@@
for all $x\in\widehat{\mathfrak{G}}^{\vee}_{\mathfrak{w}}$. More explicitly, these read as the following mode relations
@@
To prove point (3), mutual locality, we will directly compute the commutator $\mathcal{J}^{a}(z),\mathcal{J}^{b}(w)]$ between two raviolo fields and show that it is of the form in Lemma 7.1.4. That is, we will show that
(7.2.10) $\mathcal{J}^{a}(z),\mathcal{J}^{b}(w)]=(\bm{\kappa}^{-1})^{ab}\partial_{w}\Delta(z-w)+\mathbf{\tilde{f}}^{ab}_{c}\Delta(z-w)\mathcal{J}^{c}(w)\,.$
This computation involves a few grading-related technicalities, so we will present it in full detail.
To begin with, although we have been using the same Lie algebra index for both $\tilde{L}$} and $\tilde{H}$}, in fact these fields take values in two distinct vector spaces. It will therefore be important to distinguish them in what follows. We thus introduce a collective index $a=(A,\alpha)$ so that
(7.2.11) $\mathcal{J}^{a}(z)=\sum_{n\geq 0}z^{n}{\tilde{L}}^{A}_{n}+\Omega_{z}^{n}{\tilde{H}}^{\alpha}_{n}\,.$
Furthermore, recall from (6.1.14) that $\tilde{L}_{n}^{A}$ has degree $1$ while $\tilde{H}_{n}^{\alpha}$ has degree 0, and that $z^{n}$ has degree zero and $\Omega_{z}^{n}$ has degree $1$. As such, the raviolo field $\mathcal{J}^{a}(z)$ has homogeneous degree $1$ in $\mathcal{K}^{\bullet}_{z}\otimes\mathfrak{G}^{\vee}$. The commutator can thus be computed as
(7.2.12) ${\mathcal{J}}^{a}(a),{\mathcal{J}}^{b}(w)]=\sum_{n,m\geq 0}z^{n}w^{m}[{\tilde{L}}_{n}^{A},{\tilde{L}}_{m}^{B}]-z^{n}\Omega^{m}_{w}[{\tilde{L}}_{n}^{A},{\tilde{H}}_{m}^{\beta}]\\ +\Omega_{z}^{n}w^{m}[{\tilde{H}}_{n}^{\alpha},{\tilde{L}}_{m}^{B}]+\Omega_{z}^{n}\Omega_{w}^{m}[{\tilde{H}}_{n}^{\alpha},H_{m}^{\beta}]\,$,
where a minus sign appeared in the second term from commuting $\tilde{L}_{n}^{A}$ past $\Omega_{w}^{m}$, both of which have odd degree 11.121212We thank Niklas Garner for pointing this out.
At this point we must be careful when replacing the mixed commutators in (7.2.12) by their explicit expressions, which we recall here for convenience:
[…]
Recall that $(\tilde{\mu}_{2})^{\alpha B}{}_{C}$ and $(\kappa^{-1})^{\alpha B}$ were defined in (6.1.4) and (6.2.3), respectively, as the structure constants of the dual maps
(7.2.14) $\mu_{2}^{\vee}:V^{\vee}\times\mathfrak{n}^{\vee}\to\mathfrak{n}^{\vee},\qquad\langle\,\cdot\,,\,\cdot\,\rangle^{\vee}:V^{\vee}\times\mathfrak{n}^{\vee}\to{\mathbb{C}}.$
Hence, in order to write the commutator ${\tilde{L}}_{n}^{A},{\tilde{H}}_{m}^{\beta}]$ appearing in (7.2.12), we introduce the map
(7.2.15) $\nu_{2}^{\vee}:\mathfrak{n}^{\vee}\times V^{\vee}\to\mathfrak{n}^{\vee}\,,\qquad\nu_{2}^{\vee}(s^{A},t^{\beta})=-\,\mu_{2}^{\vee}(t^{\beta},s^{A})\,$,
so that, in components,
$(\tilde{\nu}_{2})^{A\beta}_{C}=-(\tilde{\mu}_{2})^{\beta A}_{C}\,.$
In complete analogy, we define
(7.2.16) $\eta^{-1}:\mathfrak{n}^{\vee}\times V^{\vee}\to\mathbb{C}\,,\qquad\eta^{-1}(s^{A},t^{\beta})=\kappa^{-1}(t^{\beta},s^{A})\,$,
such that
$(\eta^{-1})^{A\beta}=(\kappa^{-1})^{\beta A}\,.$
These definitions are chosen precisely so as to ensure the skew-symmetry of the bracket ${\tilde{L}}_{n}^{A},{\tilde{H}}_{m}^{\beta}]=-[{\tilde{H}}_{m}^{\beta},{\tilde{L}}_{n}^{A}]$. With these conventions in place, we may rewrite (7.2.12) as
(7.2.17) ${\mathcal{J}}^{a}(z),{\mathcal{J}}^{b}(w)]=\sum_{n,m\geq 0}-z^{n}\Omega_{w}^{m}\left[(\tilde{\nu}_{2})^{A\beta}_{C}{\tilde{L}}_{n-m}^{C}+m(\eta^{-1})^{A\beta}\delta_{m-1,n}\right]\\ +\Omega_{z}^{n}w^{m}\left[(\tilde{\mu}_{2})^{\alpha B}_{C}{\tilde{L}}_{m-n}^{C}+n(\kappa^{-1})^{\alpha B}\delta_{n-1,m}\right]+\Omega_{z}^{n}\Omega_{w}^{m}\tilde{f}^{\alpha\beta}_{\gamma}{\tilde{H}}_{n+m}^{\gamma}\,.$
Turning now to the right-hand side of equation (7.2.10), we compute
(7.2.18) $\Delta(z-w)\mathcal{J}^{c}(w)=\sum_{n,m\geq 0}w^{n+m}\Omega_{z}^{n}{\tilde{L}}_{m}^{C}+\Omega_{z}^{n}\Omega_{w}^{m-n}{\tilde{H}}_{m}^{\gamma}-z^{n}\Omega_{w}^{n-m}{\tilde{L}}_{m}^{C}$
which is an element of
(7.2.19) $(\mathcal{K}_{w}^{0}\otimes\mathcal{K}_{z}^{1}\otimes\mathfrak{n}^{\vee})\oplus(\mathcal{K}_{w}^{1}\otimes\mathcal{K}_{z}^{1}\otimes V^{\vee})\oplus(\mathcal{K}_{w}^{1}\otimes\mathcal{K}_{z}^{0}\otimes\mathfrak{n}^{\vee})\,.$
We define the linear map $\mathbf{\tilde{f}}^{ab}_{c}$ acting on this space as the endomorphism:
(7.2.20) $\mathbf{\tilde{f}}^{ab}_{c}=\iota_{1}\circ(\tilde{\mu}_{2})^{\alpha B}_{C}\circ p_{1}+\iota_{2}\circ\tilde{f}^{\alpha\beta}_{\gamma}\circ p_{2}+\iota_{3}\circ(\tilde{\nu}_{2})^{A\beta}_{C}\circ p_{3}$
where $\iota_{i},p_{i}$ with $i=1,2,3$ are inclusions and projections, respectively, of the direct sum decomposition (7.2.19). By construction, we then have
(7.2.21) $\mathbf{\tilde{f}}^{ab}_{c}\Delta(z-w)\mathcal{J}^{c}(w)=\sum_{n,m\geq 0}w^{n+m}\Omega_{z}^{n}(\tilde{\mu}_{2})^{\alpha B}_{C}{\tilde{L}}_{m}^{C}+\Omega_{z}^{n}\Omega_{w}^{m-n}f^{\alpha\beta}_{\gamma}{\tilde{H}}_{m}^{\gamma}\\ -z^{n}\Omega_{w}^{n-m}(\tilde{\nu}_{2})^{A\beta}_{C}{\tilde{L}}_{m}^{C}\,.$
In complete analogy, we have that
(7.2.22) $\partial_{w}\Delta(z-w)=\sum_{n\geq 0}nw^{n-1}\Omega_{z}^{n}+(n+1)z^{n}\Omega_{w}^{n+1}\,$,
which can be thought as an element of
(7.2.23) $({\mathcal{K}}_{w}^{0}\otimes{\mathcal{K}}_{z}^{1})\oplus({\mathcal{K}}_{w}^{1}\otimes{\mathcal{K}}_{z}^{0})\,$,
so that we may define the contraction with $(\bm{\kappa}^{-1})^{ab}$ as a linear endomorphism of the above space given by
(7.2.24) $(\bm{\kappa}^{-1})^{ab}=\iota_{1}\circ(\kappa^{-1})^{\alpha B}\circ p_{1}-\iota_{2}\circ(\eta^{-1})^{A\beta}\circ p_{2}\,$,
where again, $\iota_{j},p_{j}$ with $j=1,2$ are inclusions and projections, respectively, of the direct sum decomposition (7.2.23). By definition, we thus obtain
(7.2.25) $(\bm{\kappa}^{-1})^{ab}\partial_{w}\Delta(z-w)=\sum_{n\geq 0}n(\kappa^{-1})^{\alpha B}w^{n-1}\Omega_{z}^{n}-(n+1)(\eta^{-1})^{A\beta}z^{n}\Omega_{w}^{n+1}\,.$
Putting together equations (7.2.21) and (7.2.25), and comparing with (7.2.17) we finally have
(7.2.26) $\mathcal{J}^{a}(z),\mathcal{J}^{b}(w)]=(\bm{\kappa}^{-1})^{ab}\partial_{w}\Delta(z-w)+\mathbf{\tilde{f}}^{ab}_{c}\Delta(z-w)\mathcal{J}^{c}(w)\,$
as desired.
@@
Finally, point (4) follows immediately by construction of $\mathcal{V}$, and thus by the reconstruction Theorem 7.1.6, $\mathcal{V}$ defines a unique raviolo vertex algebra structure.